Case 3.2 Charlotte Honey Bee Company

Charlotte Honey Bee Company (CHBC) is considering the acquisition of equipment that would cost $200,000 to acquire and set up. CHBC expects a useful life of six years from this equipment, at which time it estimates that it could sell the equipment for $20,000. CHBC expects to generate net income of $50,000 each year before depreciation.

A. If CHBC uses straight-line depreciation, what is its expected net income after depreciation but before tax for this equipment each year?

B. If this equipment is classified as a 5-year asset for tax purposes, what is CHBC’s depreciation for tax purposes each year?

C. What, if any, is the gain or loss for tax purposes if CHBC sells the equipment at the end of six years for $20,000?

D. If CHBC’s income before depreciation is the same for both financial reporting and tax purposes, what is the difference in income for reporting and tax purposes each year? [Caution: Remember to include your result from Part C, if appropriate.]

E. What method of depreciation for financial reporting purposes do you recommend? Explain your recommendation.

To approve a single suggestion, mouse over it and click "✔"
Click the bubble to approve all of its suggestions.
to use Ginger
Limited mode
depreciation, but
×
Answer #1

a. Depreciation each year under straight line = (200,000-20,000)/6 = $30,000

Net Income each year after depreciation = 50,000 -30,000 = $20,000.

b. If it is a 5 year asset, the depreciation each year = (200,000-20,000)/5 = 36,000

c. There is no gain or loss if it sells if it sells the equipment at 20,000 since this is the salvage value.

d. Income reported for reporting would be 50,000 and that reported for tax purposes would be 20,000 for each of the first five years and in the sixth year, the income reported will be 70,000 for reporting and 40,000 for tax because of the sale.

e. I would recommend using the MACRS depreciation for these assets. The MACRS depreciation has different depreciation rates for different years which can be helpful is negating tax in the initial years of operation.